Can i Please ask for extra assistance on solving this simultaneous equation by the SUBSTITUTION Technique? Would really appreciate any help I can get.
4x+y=9
3x+2y=13
THANK YOU!

Answer by thejackal(72)

4x + y = 9
3x + 2y = 13
pick any equation, say equation 1
4x + y = 9
express y as a function of x
thus
y = 9 - 4x
plug this equation (otherwise known as substitute) into equation 2
3x + 2(9 - 4x) = 13
3x + 18 - 8x = 13
simplify
18 - 13 = 8x - 3x
5 = 5x
1 = x
plug this result back into the equation for y
y = 9 - 4x = 9 - 4(1) = 5
so y = 5 and x = 1
trying using the values of y and x in the original equations to prove that you are right